A business graduate wants to get a job in any one of the top 10 accounting firms. Applying to any of these companies requires a lot of effort and paperwork and is therefore costly. She estimates the cost of applying to each of the 10 companies and the probability of getting a job offer there. These data are tabulated below. The tabulation is in the decreasing order of cost
Company 1 2 3 4 5 6 7 8 9 10
Cost (S) 870 600 540 500 400 320 300 230 200 170
Probability 0.38 0.35 0.28 0.20 0.18 0.18 0.17 0.14 0.14 0.08
a) If the graduate applies to all 10 companies, what is the probability that she will get
atleast one offer?
b) If she applies to companies 2. 5. 8. and 9, what is the total cost?

Answer by ikleyn(52775)

.

(a)  The probability that she will NOT get any offer is

     P = (1-0.38)*(1-0.35)*(1-0.28)*(1-0.2)*(1-0.18)*(1-0.18)*(1-0.17)*(1-0.14)*(1-0.14)*(1-0.08) = 0.088149.



     The probability that she will get at least one offer is the COMPLEMENT to it

     P' = 1 - 0.088149 = 0.911851.      ANSWER




(b)  Question (b) does relate NEITHER to Algebra NOR to the Probability theory.

     It relates to Arithmetic.


     To answer this question, simply add the respective costs.